for which of the following equations is y a function of x?
a) x=y b) x=[y]

c) x=y² d) x=[y²-y]
* - [] - mean absolute value bars

Answer by stanbon(75887)

for which of the following equations is y a function of x?
a) x=y
yes
-----------------
b) x=[y]
Since (1,1) and (1,-1) are possible, it is not a function.
=================

c) x=y²
Since (1,1) and (1,-1) are possible, it is not a function.
=================
d) x=[y²-y]
x = y^2-y or x = y-y^2
not a function and not a function
The 1st is a parabola opening to the right so it has 2 y-values for
each x value in its domain except at the vertex.
The same is true for the 2nd parabola which opens to the left.
============
Cheers,
Stan H.